OpenStudy (anonymous):

What are the next three terms in the pattern 2, 6, 18, 54, ...?

Mehek (mehek14):

do you know the pattern?

OpenStudy (anonymous):

no

Mehek (mehek14):

it could be addition, subtraction, multiplication, or division by the same number so what +,-,•, or / by 2 gives you 6?

OpenStudy (anonymous):

i got it

Mehek (mehek14):

what did u get?

OpenStudy (anonymous):

162

Mehek (mehek14):

correct :) the other terms?

OpenStudy (anonymous):

they all multiply by 3

Mehek (mehek14):

yup

Mehek (mehek14):

now you can get the next 3 terms \(\ddot\smile\)

OpenStudy (anonymous):

\[2*3=6,6*3=18,18*3=54,.......... \]

OpenStudy (crabbyoldgamer):

You have found one possible set of "next three terms" in the pattern. A sequence can have multiple rules. The rule you found was "the next number is the previous number multiplied by 3". Another rule could be "the first six numbers are the previous number multiplied by 3, while the remaining numbers are all 10000". Your answer is correct. But keep in mind, a sequence can have multiple rules.

OpenStudy (anonymous):

\[2\left( 3^0,3^1,3^2,3^3,3^4,3^5,3^6,... \right)\]